OpenStudy (hanna1):

will medal and fan!! Which ordered pairs, in the form (x,y), are solutions to the equations? 3x-y/4=11 (3,8), (4, 4)

KjSaif (kjsaif):

3x-(y/4)=11 subsitute in (x,y) form if (3,-8) subsitute 3 for x and -8 for y 3(3)-(-8/4)=11 9-(-2)=11 9+2=11 11=11 (3,-8) is a valid soultion try 4,4 subsittue 4 for x and 4 for y 3(4)-(4/4)=11 12-(1)=11 12-1=11 11=11 (4,4) is a valid solution both are true so the answer,so.... both are solutions :) ~~~~Apologiabiology